# DiffDecompose: Layer-Wise Decomposition of Alpha-Composited Images via Diffusion Transformers

### v2

@@ Appendix A AlphaBlend dataset
As can be seen from Figure  8, the properties of the AlphaBlend dataset are different from the existing datasets that are intended to generate a foreground layer object information that is completely opaque but has an alpha channel. The AlphaBlend datasets consider different semi-transparent conditions, from layer-level color distanglement to simple region-level transparent object occlusion. The task is no longer the same as the tasks of object removal [69, 39, 17], downstream image restoration [29, 32], or layer generation [71, 64], and is also a common scenario in the real world. For tasks I and II, the rain, fog, light, and dirt can always occlude the whole background. Different from the common image dehazing and deraining, the pixel of the background has not been covered or fuzzy, but makes itself present semi-transparent. For task III, we further consider the more complex watermark condition in the real world, the current watermark dataset always presents lower opacity, and smaller occlusion, providing sufficient background information for inpainting methods to predict their background. Thus, we introduce bigger coverage with the size of 96$\sim$128, and the opacity below 0.25, which is more difficult to recognize. Tasks III, V, and VI, respectively, represent different pixel-level non-linear occlusions. These subtasks utilize part of the dataset [75, 1, 57, 50].
@@ Appendix B Training Loss Function
The loss function is formulated within the conditional flow matching [19] framework as the joint expectation over timesteps, noisy latent variables, and noise samples:
@@ Appendix D More Comparison Experiments and Generative Results
We use the Frechet Inception Distance (FID) [23] to assess the photorealism of the generated images by comparing the source image distribution with the inpainted image distributions. To measure the accuracy of correct object removal, we introduce root mean squared error (RMSE)[24], structural similarity index measurement (SSIM)[56], and Learned Perceptual Image Patch Similarity (LPIPS) [45].
@@ D.2 More Comparisons
For our proposed AlphaBlend, since the current inpainting methods [69, 39, 17] mainly rely on the mask-based method to achieve image inpainting, the task of occlusion and light removal is not reasonable for them to edit the whole image as shown in Figure 9, no matter how we tune the strengthen and change the prompt, it is still difficult to remove the rain and light. Thus, we only compare the results of the cell, glassware, and watermark, which can also be regarded as the region-level task instead of the layer-level task for these competing methods. It can be seen that when facing the transparent scenarios, the current methods directly predict the mask region with the background or prompt instead of fully utilizing the information of its transparent region. Additionally, when the occlusion region is pixel-additive, the competing method is prone to utilize the background to generate an unreasonable object to fix the removal cell. These results suggest that when the target object is no longer a simple pixel-level coverage, introducing the region-based methods for image editing is difficult to recover accurately. In comparison, our proposed DiffDecompose can better remove the foreground layer and maintain the background layer successfully.

### v3

@@ 5 Conclusion
Overview In the Appendix, we first present more details about our proposed dataset, AlphaBlend. Then, we introduce more information about the training and inference of DiffDecompose. Subsequently, we offer more competing results upon our proposed AlphaBlend dataset and a public dataset, LOGO, and also include additional quantitative experiments and complexity analyses. Finally, we present more visual results from generative results of DiffDecompose. The supplementary includes the following sections:
  • A.AlphaBlend dataset.
  • B.Training Loss Function.
  • C.Inference Algorithm.
  • D.More Comparison Experiments and Generative Results.
  • E.User Study and Broader Impacts.
  • F.Limitations.
@@ Appendix A AlphaBlend dataset
[…]
@@ Appendix B Training Loss Function
The loss function is formulated within the conditional flow matching flowmatch framework as the joint expectation over timesteps, noisy latent variables, and noise samples:
@@ Appendix D More Comparison Experiments and Generative Results
We use the Frechet Inception Distance (FID) heusel2017gans to assess the photorealism of the generated images by comparing the source image distribution with the inpainted image distributions. To measure the accuracy of correct object removal, we introduce root mean squared error (RMSE)hodson2022root, structural similarity index measurement (SSIM)wang2002universal, and Learned Perceptual Image Patch Similarity (LPIPS) snell2017learning.
@@ D.2 More Comparisons
For our proposed AlphaBlend, since the current inpainting methods yu2023inpaint; sdxl; ekin2024clip mainly rely on the mask-based method to achieve image inpainting, the task of occlusion and light removal is not reasonable for them to edit the whole image as shown in Figure 9, no matter how we tune the strengthen and change the prompt, it is still difficult to remove the rain and light. Thus, we only compare the results of the cell, glassware, and watermark, which can also be regarded as the region-level task instead of the layer-level task for these competing methods. It can be seen that when facing the transparent scenarios, the current methods directly predict the mask region with the background or prompt instead of fully utilizing the information of its transparent region. Additionally, when the occlusion region is pixel-additive, the competing method is prone to utilize the background to generate an unreasonable object to fix the removal cell. These results suggest that when the target object is no longer a simple pixel-level coverage, introducing the region-based methods for image editing is difficult to recover accurately. In comparison, our proposed DiffDecompose can better remove the foreground layer and maintain the background layer successfully.